Julia is five times as old as her son. Six years from now, Julia will be three times as old as her son. What's Julia's age?
Answer by TimothyLamb(4379)

j = 5s
j + 6 = 3(s + 6)
j + 6 = 3s + 18
---
j + 6 = 3s + 18
5s + 6 = 3s + 18
2s = 12
---
s = 6
j = 30
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php